### v3

# Robotic Sorting Systems: Robot Management and Layout Design Optimization

Abstract
In the contemporary logistics industry, automation plays a pivotal role in enhancing production efficiency and expanding industrial scale. In particular, autonomous mobile robots have become integral to modernization efforts in warehouses. One noteworthy application in robotic warehousing is the robotic sorting system (RSS), which is distinguished by its cost-effectiveness, simplicity, scalability, and adaptable throughput control. Previous research on RSS efficiency often assumed an ideal robot management system, ignoring potential traffic delays and assuming constant travel times. We introduce a novel robot traffic management method, named Rhythmic Control for Sorting Scenario (RC-S), for RSS operations, along with an estimation formula that establishes the quantitative relationship between system performance and configurations. Simulations validate that RC-S reduces average service time by 10.3% compared to the classical cooperative A* algorithm, while also improving throughput and runtime. Based on the performance analysis of RC-S, we develop a layout optimization model that considers system configurations, desired throughput, and costs to minimize expenses and determine the optimal layout. Numerical studies show that facility costs dominate at lower throughput levels, while labor costs prevail at higher throughput levels. Additionally, due to traffic efficiency limitations, RSS is well-suited for small-scale operations like end-of-supply-chain distribution centers.
Keywords: Logistics; Robotic sorting system; Robot management system; Performance evaluation; Layout design
1 Introduction
Improved logistics and delivery services have fueled the rapid growth of e-commerce in the 21st century. The development of stable supply chains, same-day or next-day delivery options, and hassle-free return policies increase consumer confidence in online shopping (BusinessWire,, 2023). In order to handle the increasing volume of online orders, a new generation of warehouses specifically catering to individual customers has become a hot topic for logistics companies. This kind of warehouse efficiently meets the demand for small orders with tight delivery schedules through the implementation of automated equipments (Boysen et al., 2019b, ). For instance, robotic mobile fulfill system (RMFS), which employs a rack-moving mechanism, is widely used in intra-warehouse logistics, such as Amazon’s KIVA systems (Wurman et al.,, 2008).
The main activities performed in warehouse include: (1) receiving, (2) transfer and put away, (3) order picking/selection, (4) accumulation/sorting, (5) cross-docking, and (6) shipping (De Koster et al.,, 2007). In this paper, we focus on the sorting process. Sorting involves categorizing and consolidating parcels according to their order information and shipping destinations. Conventional sorting systems commonly employ conveyor-based sorters , where the actuators move along with the conveyor belt, sequentially passing through each outbound station and releasing the loaded parcels at the appropriate locations (Boysen et al., 2019a, ). These systems are highly appreciated for its efficiency and stability, while suffering from the inflexibility and significant space occupation (Boysen et al.,, 2023).
(a) A dual-layer RSS solution
(b) Conveyor-belt sorting robot
Figure 1: The application of RSS (source: www.geekplus.com)
A novel sorting system utilizing mobile robots attracts increasing attention in recent years, which is called the robotic sorting system (RSS) (Zou et al.,, 2021). Figure 1(a) presents a dual-layer solution of RSS. The site is arranged with numerous delivery ports, each associated with a specific sorting category, gathering all the parcels of that category. A delivery port can be a bin, a cage cart, or a chute in a dual-layer structure. Robots deliver parcels to delivery ports, serving as the combination of conveyor belts and actuators in conventional sorting systems. Loading stations are located on the periphery of the site, where workers load parcels onto the robots. A top-mounted tray or conveyor actuator enables the robot to load or unload parcels. Figure 1(b) shows a specific model of sorting robot developed by Geekplus, a global technology company specialized in smart logistics and robots. The process for a robot to execute a sorting task is as follows:
  1. 1.
    The robot receives a parcel at a loading station, along with the destination information;
  2. 2.
    The robot moves along aisles to the designated outlet;
  3. 3.
    When reaching the target outlet, the robot releases the loaded parcel;
  4. 4.
    The robot returns to the waiting zone behind a loading station and becomes idle.
E-commerce order fluctuate during special selling seasons; therefore, warehouse throughput must be designed to accommodate these variations. Unlike conveyor-based sorting systems, the independent and modular nature of robotic operations contributes to the flexibility and scalability of RSS, making it well-suited to dynamic demands of e-commerce (Azadeh et al.,, 2017). During off-hours, warehouse managers can easily configure the control software to change the status of loading stations, the correspondences between outlets and destinations, and the aisle network topology (Xu et al.,, 2022). To leverage the flexibility, robotics companies introduce the innovative Robots-as-a-Service (RaaS) business model, allowing logistics clients to rent robots as needed. RaaS eliminates the high costs associated with purchasing and maintaining robots and peripheral equipment, enabling warehouse managers to adjust the number of rented robots based on demand fluctuations.
Despite these advantages, a research gap remains in investigating the flexibility of RSS and minimizing system costs, as accurately modeling system efficiency is challenging. In robotic systems, robots rely on path planning and coordination algorithms to complete tasks. The problem of finding conflict-free trajectories for all robots is known as the multi-agent path finding (MAPF) problem. The MAPF problem has been proven to be NP-hard (Yu and LaValle,, 2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in the instability of RSS efficiency. Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in RSS, particularly when dealing with a substantial parcel flow. Zou et al., (2021) proposed several closed queuing network (CQN) models to quantitatively analyze the impact of congestion and designed an algorithm to estimate throughput. Validated through numerical experiments, these CQN models accurately reflect the performance of a real case from Deppon Express. However, the models primarily focus on the process of agents transferring between queues rather than on traffic flow and do not fully consider traffic management methods to mitigate conflicts. This aspect makes them less suited for analyzing the efficiency of MAPF solvers, such as the method proposed in this paper. On the other hand, other existing studies often overlooked issues related to robot coordination, relying on the idealized assumption that there are no conflicts or deadlocks. Consequently, their results may become distorted as the actual travel time of robots deviates from the free-flow travel time. Furthermore, these studies do not adequately explore the flexibility of RSS, particularly the ability to adjust the number of robots.
Given the existing gap, this paper aims to address critical challenges in designing an efficient RSS, focusing on three key areas. First, we propose a novel robot traffic management method to enhance system throughput, ensuring stable and efficient operations in multi-robot collaboration settings while overcoming the complexities of robot interactions. Second, we develop a precise model to quantify system efficiency in RSS, enabling accurate throughput estimation to clarify the relationship between system configurations and sorting capacity. Third, we highlight the need for management strategies that leverage RSS flexibility to provide cost-effective responses to fluctuating demands. These contributions aim to improve the performance and scalability of RSS implementations.
To address the aforementioned challenges, we first explore the potential of centralized control strategy for sorting robots. To fully utilize global information while considering robot operational patterns and safety distance regulations, we adapt a network-level control strategy for coordinating autonomous vehicles and propose a collision-free spatio-temporal path planning method for RSS. An efficient heuristic algorithm is proposed to enhance the scalability. Second, building upon this control strategy, we derive an estimation formula for throughput in RSS given system configurations, which serves as an efficiency constraint in the planning problem. Finally, we propose an optimization model aimed at minimizing the initial investment and average operations costs under fluctuating demands. The solution includes layout design and resource allocation recommendations, specifically on how to adjust the number of robots and workers. Experiments provide insights into the cost structure of RSS under varying unit price conditions. To the best of our knowledge, this paper is among the first to incorporate the traffic issues of robots during operations into the configuration design of the RSS system. Our intention is to bridge operational and strategic planning in multi-robot coordination contexts, thereby contributing a comprehensive theoretical framework to support decision-making in such complex and flexible systems.
The remainder of this paper is organized as follows. In section 2, we provide an extensive review of relevant research on RSS systems and conventional sorting systems, highlighting the contributions of this research. Section 3 offers a detailed problem description. In section 4, we introduce an innovative traffic management framework to coordinate multiple robots in a warehouse setting. Based on the proposed framework, section 5 presents a system throughput estimation formula, which provides quantitative measurement of layout configuration’s impact to system efficiency. In section 6, we demonstrate the superiority of our proposed traffic management framework compared with benchmarks, as well as validate the accuracy of throughput estimation formula. To emphasize the impact of traffic issues on RSS, experiments are conducted to show that queueing network model has estimation biases for throughput in certain scenarios. Moving on to section 7, we propose a layout optimization model to minimize the total costs and introduce an efficient solution approach. Section 8 presents a sensitivity analysis to investigate the optimal layout design and the corresponding cost structure under different unit cost, and insights distilled from the results. Finally, section 9 provides a summary of the entire paper.
2 Literature Review
The cost savings in labor through automation are particularly crucial in the picking process, and increased sorting efficiency enables parcel companies to offer convenient and low-cost same-day delivery services (Dekhne et al.,, 2019). However, the operations research community has scarcely explored the realm of robotized sorting systems, which represent the latest advance of warehouse automation. Among the existing literature, studies mainly focus on the efficiency of small parcel sorting systems in distribution centers. This body of work delves into the comprehensive assessment of transportation efficiency within the robot fleet under various system configurations. In the majority of these studies, robots’ sorting tasks are modeled as services in queuing network models, and the total throughput is the key performance evaluation metrics (Zou and Chen,, 2020; Zi and Gao, 2020a, ; Zi and Gao, 2020b, ; Zou et al.,, 2021; Xu et al.,, 2022). Zi and Gao, 2020a ; Zi and Gao, 2020b assumed that each robot moves independently. Zou and Chen, (2020) introduced a semi-open queuing network (SOQN) model and subsequently improved its theoretical aspects in their follow-up research (Zou et al.,, 2021), simplifying it to the closed queueing model (CQN) by assuming a sufficient high arrival rate of parcels. A queueing model takes into account the queuing behavior of robots at loading stations and outlets, partially addressing the issue of robot occupancy on the aisle. Furthermore, their second study compares how the minimum total costs of the system varied under different network topology and different unit cost of system components, given the target sorting efficiency. Xu et al., (2022) further investigated an RSS with a parcel-to-loading-station assignment mechanism, which could be regarded as a pre-sorting strategy. This study removes the constraint on the number of robots, resulting in an open queuing network (OQN) model. Experimental results demonstrates that the introduction of pre-sorting effectively reduces the average travel distance of the robots but might lead to extra congestion upstream at the same time.
Liu et al., 2019b and Tan et al., (2021) focused on the task assignment problem with the objective of minimizing the sorting makespan. Both adopts a travel time model (closed-form travel time expression, defined by Azadeh et al., (2017)), and the second research formulates a mixed-integer programming model. Compared to queuing network models, travel time models exhibit significant distortion when the number of robots is high, as they allow multiple robots to simultaneously occupy a facility. Boysen et al., (2023) conducted research on order sorting system, which involves additional tasks including Piece-to-Order assignment and Order-to-Collection Point assignment. Therefore, the primary focus of the referred study is not solely on scheduling and control of robots, and the travel time of robots to a specific outlet is assumed to be constant. Shi et al., (2021)’s study, on the other hand, investigates a human-robot hybrid sorting system, in which the scale of robot sorting is fixed, while the manual capacity is adjustable according to the demand. Regrettably, in this context, the considerable flexibility inherent in the robotic system has not been fully harnessed.
Table 1: Literature summary on RSS investigation
Reference System Modeling Objective Congestion & Deadlock
Liu et al., 2019b parcel sortation TT, MIP MS, C $\times$
[…]
Zi and Gao, 2020a parcel sortation Q T $\times$
Zi and Gao, 2020b parcel sortation Q, MIP T, C $\times$
[…]
Tan et al., (2021) parcel sortation TT, MIP MS $\times$
Xu et al., (2022) parcel sortation OQN T $\times$
Boysen et al., (2023) order sortation TT, MIP T $\times$
This paper parcel sortation TT, MIP T, TC $\checkmark$
[…]
Throughout the existing research, a limited number of studies consider the floor space cost while others relax the constraints on aisle resources. Furthermore, finding collision-free paths for multiple vehicles has been proved to be an NP-hard problem (Surynek,, 2010), and some state-of-the-art algorithms also struggle to completely eliminate the issue of road congestion resulting from an increasing number of robots. However, except the study by Shi et al., (2021), which employs a fitting method to predict the positive correlation between traffic flow and congestion, most of the studies neglect the mutual influence among multiple robots in aisle occupation for a compact system, leading to an misestimate of traffic capacity. Table 1 presents a comprehensive overview of the literature reviewed on the RSS, highlighting its relevance to our paper.
Some researchers recognized the limitations of queuing models in capturing the impact of aisle layout design on multi-robot systems and chose to address the traffic issues by modeling them as MAPF problems. Wagner and Choset, (2011) proposed a novel M* algorithm that allowed for the planning of a larger number of robot trajectories. However, they faced challenges due to significant memory requirements and time complexity, and could not guarantee a feasible solution. Some researchers designed priority-based algorithms and integrated task assignment to reduce waiting time (Nguyen et al.,, 2019; Liu et al., 2019a, ), sacrificing solution optimality in exchange for stability and computational efficiency. Li et al., (2021) introduced an RHCR algorithm framework to tackle the windowed MAPF problem, providing an efficient solution for managing newly arrived orders in warehouse operations. However, despite their effectiveness, MAPF-based studies have faced challenges in deriving a closed-form expression for throughput estimation, which is crucial for creating reliable and efficient system configurations. Consequently, few studies on multi-robot systems comprehensively address both operational robot management and strategic layout design while considering their interplay.
In summary, existing research on RSS systems lacks sufficient focus on robot traffic management and overall system costs. This study addresses these gaps by first developing a dedicated robot traffic management method for RSS at the operational level. Building on this method, we establish an optimal layout and configuration design model aimed at minimizing overall system costs in the long run, ensuring that the results are practically applicable in real-world warehouses.
3 Problem Description
In an RSS, three zones are deployed from center to periphery in the field (Wang et al.,, 2021): 1) sorting zone; 2) loading zone; 3) waiting zone, as shown in Figure 2. Parcels are loaded to robots by workers in the loading zone. Robots drop off each parcel to its target outlet in the sorting zone to fulfill a sorting task. Waiting zone is the space where robots idle and queue behind loading stations. The area of every zone is determined in the design phase, and remains unchanged regardless of sorting demands, number of robots and workers. Let $W_{l}$ and $W_{w}$ denote the width of loading zone and waiting zone, respectively, which are considered to be constant in this study. The decision of site planning mainly focuses on the design of sorting zone, particularly the configuration of the aisle network.
Figure 2: Layout of RSS
Without loss of generality, we make the following assumptions in this study:
  • The parcels arrive in batch manner. When a parcel is loaded on a robot, the next parcel will arrive at the loading station immediately. That is, the loading time is negligible, and this parameter is considered implicitly in the loading rate of worker.
  • The arrival rate of parcels destinating for each outlet at each loading station is the same (When the sorting demand is unevenly distributed, this assumption can be achieved by increasing or decreasing the number of outlets corresponding to each destination. Additionally, outlets with extremely low and high arrival rates can be placed together to balance the flow);
  • The time for robots dropping parcels into outlets is negligible (To satisfy this assumption, robots need to release goods in advance based on its speed before reaching the target outlet, similar to the behavior of conveyor-based sorters. If it is difficult to achieve due to technical and cost issues, a 2-second drop-off time per delivery must be introduced (Zou et al.,, 2021));
  • The drop-off operation by robots around one outlet does not affect each other, i.e., each outlet allows at most four robots releasing parcels simultaneously;
  • Each robot can only carry one parcel during one delivery process;
  • Loading stations are evenly distributed in the loading zone;
  • The transportation of parcels after being collected in the outlet is not considered in this research.
  • The battery charging needs of robots are not considered.
(a) Sample gird layout of a compact $4+4$ aisle network
(b) Graph representation of sample layout
Figure 3: Illustration of aisle network in sorting zone
To better evaluate the performance of this system with high loads and complicated sorting demands, the microscopic traffic management of robots in the sorting zone needs careful investigation. Notably, we use the term “aisle” instead of “road” in the following content to clarify that each road in the network is single-lane. A sample grid-based map in RSS is depicted in Figure 3(a). It consists of $n_{v}$ vertical aisles and $n_{h}$ horizontal aisles. Each cell in the map has a side length of $D$. Robots locate themselves using QR codes at the center of each cell and can only move between adjacent cells. We assume that the length of sorting robots is less than $D$, allowing them to rotate within a cell. We use the term “outlet” to refer to the special cells designated for collecting parcels delivered by robots. Outlets are evenly distributed throughout the sorting zone, divided by aisles. This layout allows robots to deliver parcels to outlets on both sides of the aisle and ensures that the outlets are arranged with sufficient density. Each loading station in the loading zone is connected to a pair of dedicated entrance and exit, which are adjacent to each other. Therefore, the number of loading stations cannot exceed the the number of aisles. Figure 3(b) shows the graph representation of the sample layout in Figure 3(a). In the graph, each cell that makes up the aisles is modeled as either a conflict node (blue dot), unloading node (red hexagon), or entrance/exit node (green/yellow square), with edges of length $D$ connecting adjacent nodes. Conflict nodes are at the intersections of aisles, unloading nodes are where robots deliver parcels, and entrance/exit nodes are at aisle ends.
Figure 4: Common types of conflicts: (a) a swapping conflict, (b) a vertex conflict, (c) a following conflict, and (d) a deadlock conflict
The sorting robots share the same kinematic parameters, with a maximum velocity of $v_{max}$ and maximum acceleration/deceleration of $c_{max}$. Robots possess the capability of in-place rotation, with angular velocity denoted as $\omega_{r}$. To ensure collision-free movement of robots, the robot management system needs to avoid conflicts in path finding. Common conflicts, as illustrated in Figure 4, include the following four types (Stern et al.,, 2019): (a) Swapping conflict occurs when two robots are planned to swap locations at the same time. It does not exist in a one-way network; (b) Vertex conflict occurs when more than two robots are planned to occupy the same vertex; (c) Following conflict occurs when one robot is planned to occupy a vertex that was occupied by another robot in the previous timestep. The latter is stationary or moving at a lower speed; (d) Deadlock conflict occurs among multiple robots, with each robot having a following conflict with another robot. Among these conflicts, both swapping conflicts and deadlock conflicts necessitate a re-planning of paths. The occurrence of either can result in significant losses in system efficiency.
Figure 5: Overview of research methodology
Planing problems related to RSS can be divided into two stages: the site planning stage and operations stage. In the site planning stage, warehouse manager concentrates on the division of zones for different functions in the warehouse. The area of an RSS should be large enough to fulfill the designated throughput level in peak season, and the number of outlets should be no fewer than the number of individual sorting destinations. The decision variables of site planning stage are the number of horizontal and vertical aisles $n_{h},n_{v}$ and the number of loading stations $n_{l}$; in the operations stage, warehouse manager should adaptively adjust the number of workers and robots to minimize operations costs, given expected sorting demands. Since a warehouse is usually rented for years, it is advisable to keep the zone division unchanged during its operation. The decision variables of operations stage are the number of workers and robots in operating period $\sigma$, noted as $n_{w}^{\sigma},n_{r}^{\sigma}$, respectively.
The goal of RSS planning problem is to minimize the total costs, which consists of two parts: the facility costs $C_{f}$, and the operations costs $C_{o}$. The size of the site holds a trade-off between $C_{f}$ and $C_{o}$: A small network leads to lower capacity, thereby increasing the waiting time of robots and raising the rental cost of robots in operations costs; whereas, an overly large network leads to increased site costs and longer travel distance. An overview of research methodology is shown in Figure 5. In sections 4-6, we present the traffic management of robots, conduct system efficiency analysis, and validate the performance and accuracy of estimation formula, respectively. In section 7, a layout design optimization model is formulated to minimize the total costs of RSS given the demand.
4 Management of High-density Robot Traffic in RSS
Managing a large number of robots within a dense network poses significant challenges, especially when balancing system throughput performance with the computational load of algorithms under unpredictable traffic conditions. To address multi-robot path finding and eliminate traffic gridlock within an admissible computational duration, this study adopts an innovative autonomous vehicle management scheme, called Rhythmic Control (RC) (Lin et al.,, 2021). It enables the planning of uninterrupted scheduling for robot fleets. In this section, we begin with the key concepts inherited from the original RC: the virtual platoon and the cycle. By leveraging the reservation mechanism, we propose a new centralized framework called Rhythmic Control in Sorting Scenario (RC-S). RC-S presents high orderliness and serves as the foundation for theoretical analysis of system efficiency, which will be discussed in detail in section 5.
4.1 Virtual platoon and cycle
Lin et al., (2021) investigated a method of incorporating rhythm into traffic management at the macro-traffic level and introduced the concept of virtual platoon (VP). VPs represent the spatiotemporal slot that is generated in a rhythmic manner and keep uniform linear motion. As shown in Figure 6(a), VPs move along a designated spatiotemporal trajectory and maintain a constant safe distance between each other. Figure 6(b) shows the spatiotemporal trajectory of VPs in a four-lane network, without any collision or deceleration. We discretize time based on the generation time of VPs at the entrance, and each time interval is noted as a cycle. The duration of each time interval is the cycle length. By staggering the entry times of VPs at different entrances within the same cycle, VPs pass one intersection alternately with relatively short headway and the intersection capacity could be maximized.
Whenever a vehicle needs to pass through a lane, it is mandated to follow the movement of one VP in that lane until it reaches its destination and exits the lane. Before entering the network, the vehicle must wait for an unoccupied VP at the entrance. It first sends its destination to the control system, which then plans a conflict-free route and reserves one unit of capacity from the corresponding VPs in different segments of the route. The capacity is released when the vehicle leaves each VP. Through the reservation mechanism, RC ensures that the number of vehicles within each VP does not exceed its predetermined capacity. Consequently, the trajectories of vehicles are conflict-free.
(a) Virtual platoons in a four-lane network
(b) Trajectory of virtual platoons in different directions. (Lin et al.,, 2021)
Figure 6: Concept of Rhythmic Control
RC is essentially a traffic control method, similar to a combination of globally coordinated traffic signal control and road network flow control. Within this framework, it is essential to clarify the movement patterns of each entity for integration into a sorting-specific robot management system. For detailed information on the RC framework, we refer readers to Lin et al., (2021). The following points provide a foundational overview critical to our study on the implementation of the RC-S scheme:
  • A VP represents a time–space slot that is available for being occupied by vehicles and is generated in a rhythmic manner in each entrance of the network.
  • VPs move in a straight line at a constant speed while maintaining a fixed distance, ensuring that the total number of VPs present in the road network remains the same at any given moment.
  • After entering the network, vehicles need to continuously move along with a specific VP or transfer from one VP to another until exiting through an exit.
  • At any given moment, the number of vehicles accommodated by a VP cannot exceed its capacity. If the required VP capacity for vehicle travel is insufficient, the vehicle needs to queue and wait at the entrance until entering the network in a subsequent cycle.
4.2 Concept of RC-S
In order to embody RC in the RSS, we regulate the behavior of robots, and propose the RC-S scheme, which can be regarded as a centralized MAPF solver. To focus on the behavior of each individual robot, we discretize the map into grids, defining each grid where robots can pass through as a node. Given $(n_{v},n_{h})$, the network in sorting zone is represented by a directed graph $\mathcal{G}=(\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{e},\mathcal{E})$, where $\mathcal{V}_{c}$, $\mathcal{V}_{u}$, $\mathcal{V}_{e}$ and $\mathcal{E}$ stand for the set of conflict nodes, unloading nodes, entrance/exit nodes and edges, respectively. To reduce intersections, we adopt a one-way network, where aisles from different directions are arranged in an interleaved pattern, as shown in Figure 7. Based on the network structure, each VP can accommodate at most one AGV. Let $\tau_{e}$ and $\tau_{c}$ denote the fixed travel time of VPs on each unit cell in the grid map and the cycle length of rhythmic control, respectively. Under RC-S, the behaviors of robots are limited to three actions: moving straight, dropping off parcel, and turning. Robots enter the network with fixed interval $\tau_{c}$, thus the upper bound of robot flow on each grid is $1/\tau_{c}$.
Figure 7: Illustration of RC-S. (a)-(d) are the four phases in one cycle.
In each cycle, the distribution of VPs could be divided into four phases (Figure 7(a)-(d)), and the duration of each phase is $\tau_{e}$. We have:
$\tau_{c}=4\tau_{e}$ (1)
$\tau_{e}$ and $\tau_{c}$ are determined by the maximum velocity of robot $v_{max}$ and the maximum loading rate of each loading station $r_{l}$, respectively:
$\tau_{e}\geq\frac{D}{v_{max}}$ (2)
$\tau_{c}\geq\frac{1}{r_{l}}$ (3)
where $D$ is the length of each grid. At phase(a) of a cycle, new VPs are released at the entrance of each aisle; at phase(c), VPs leave sorting zone through the exit of each aisle. Thus, if we do not count the VPs at the exit of each aisle, then the number of VPs in the network remains the same throughout the entire sorting process. Intersections are occupied by VPs at phases(a) and (c); each outlet could be served by four VPs in four different directions at phases(b) or (d). The fixed number of VPs and the regularity in their phases allow us to manage and analyze this system from a global perspective. The speed of each VP could be derived by:
$v_{VP}$ $=\frac{D}{\tau_{e}}$ (4)
Figure 8: Illustration of turning and transfer process for one robot (red dot)
RSS does not require auxiliary turning lanes since robots are capable to rotate within the intersection grid. The process of turning and transfer between VPs for robots are demonstrated in Figure 8. As shown in the figure, the robot to turn first decelerates to zero, stops following its original VP, rotates 90 degrees at the intersection, joins a new VP and follows the movement of it. Turning action entails one cycle for both the original and subsequent VPs, leading to a reduction in traffic capacity.
During this process, the parameters of RC-S needs to satisfy the following conditions:
$d(\tau_{e},v_{VP},v_{max},c_{max})\geq 2D$ (5)
$\tau_{e}\geq\frac{\pi}{2\omega_{r}}$ (6)
Constraint (5) implies that the robot can catch up with the VP in a turn, where $d(\tau_{e},v_{VP},v_{max},c_{max})$ denotes the maximum travel distance a robot can cover in acceleration/deceleration process. The expression is derived in detail in Appendix B; Constraint (6) imposes limits on the rotation speed $\omega_{r}$, ensuring that the robot can complete a 90-degree turn within one phase. In summary, using expressions (1)-(6) and the kinematic parameters of the robots, we could determine the parameters for RC-S.
The trajectory of robot in one delivery task should start from one loading station, pass the target outlet corresponding to the destination of parcel, and end at another loading station. We define feasible path in RC-S:
DEFINITION 1.
A feasible path in RC-S is a sequence of VPs and their occupying cycles. The trajectories traveled by these VPs during the occupied cycles are connected end-to-end, linking two active loading stations at the periphery of the network.
Definition 1 defines the candidate spatio-temporal paths for robots under RC-S, where an active loading station indicates that a worker is currently engaged in tasks. When a robot performs a sorting task, it moves from one loading station $i$ to another loading station $j$, $i,j\in\mathcal{L}$, where $\mathcal{L}$ is the set of all loading stations. Let $\mathcal{O}$ represent the set of outlets. Then, we denote the set of all feasible path as $\mathcal{R}=\cup_{i\in\mathcal{L},j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(i,k,j)$, where $\mathcal{R}(i,k,j)$ is the subset containing the paths start from loading station $i$, pass by outlet $k$ and end at loading station $j$. To simplify the notation, we use $\mathcal{R}_{i,k}$ to represent the set of all feasible paths originating from station $i$ and passing by outlet $k$, and ^$\hat{\mathcal{R}}_{i}$ to represent the set of all feasible paths heading for station $i$, that is, $\mathcal{R}_{i,k}=\cup_{j\in\mathcal{L}}\mathcal{R}(i,k,j)$ and ^$\hat{\mathcal{R}}_{i}=\cup_{j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(j,k,i)$. In each cycle, the feasible path assignment problem (FPA) for all the loaded robots is formulated as follows.
$\underset{x_{i,k}^{r},\hat{x}_{i,k}}{\min}\ \sum_{i\in\mathcal{L}}\sum_{k\in\mathcal{O}}(\sum_{r\in\mathcal{R}_{i,k}}c_{i,k}^{r}x_{i,k}^{r}+\hat{c}_{i,k}\hat{x}_{i,k})$
$s.t.\quad$ $\sum_{i\in\mathcal{L}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}\delta_{i,k}^{r,\nu,l}x_{i,k}^{r}\leq N_{\nu}^{l}$ $\forall\nu\in\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{e},l\in\mathcal{C}$ (7)
$\hat{x}_{i,k}+\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=d_{i,k}$ $\forall i\in\mathcal{L},k\in\mathcal{O}$ (8)
$\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=\sum_{j\in\mathcal{L}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{j,k}\cap\hat{\mathcal{R}}_{i}}x_{j,k}^{r}$ $\forall i\in\mathcal{L}$ (9)
$x_{i,k}^{r},\hat{x}_{i,k}\in\{0,1\}$ $\forall i\in\mathcal{L},k\in\mathcal{O},r\in\mathcal{R}_{i,k}$ (10)
There are two sets of variables: $x_{i,k}^{r}$ equals one if the feasible path $r\in\mathcal{R}_{i,k}$ is reserved in current cycle by a robot, and zero otherwise; $\hat{x}_{i,k}$ equals one if the robot at loading station $i$ is heading for outlet $k$ and is kept waiting until the next cycle, and zero otherwise. In the objective function, $c_{i,k}^{r}$ is the travel time on path $r\in\mathcal{R}_{i,k}$, and $\hat{c}_{i,k}$ is the penalty of delay. The objective function is to minimize the total costs in this cycle. Constraint (7) indicates that in each cycle, at most one robot can be present at node $\nu$, where $N_{\nu}^{l}$ is the updated remaining capacity in node $\nu$ in cycle $l$, $N_{\nu}^{l}\leq 1$. Parameter $\delta_{i,k}^{r,\nu,l}$ equals one if feasible path $r$ will occupy node $\nu$ in cycle $l$ and zero otherwise. Constraint (8) describes the path allocation results for a robot at loading station $i$, where $d_{i,k}$ equals one if the robot is heading for outlet $k$, and zero otherwise. Constraint (9) is the conservation of flow constraint, ensuring a stable number of robots at each loading station.
In large-scale network, the number of feasible paths rapidly increases. It is difficult to solve FPA within a reasonable amount of time. Here we propose a heuristic algorithm to generate a feasible solution in a relatively short time, as shown in Algorithm 1. To expand the search depth, feasible paths starting at some future cycle are allowed to be reserved. We further define the entry cycle:
DEFINITION 2.
The entry cycle of a specific feasible path is the earliest cycle at which all VPs in the path are unoccupied.
Algorithm 1 A heuristic method to implement RC-S
0:  Load the reservation table that records the status of VPs on the map within $N_{p}$ cycles. Set the maximum search range in each cycle as $N_{t}$. Calculate feasible path set $\mathcal{R}$.
0:  At the beginning of cycle $t$, add the first robot of each queue at entrance to the candidate list.
0:  If the candidate list is empty, go to step 5; else, find the robot with the longest waiting time in the candidate list, obtain its origin $i$ and target outlet $k$. Calculate the entry cycle of each feasible path in $\mathcal{R}_{i,k}$, and choose the feasible path with the smallest entry cycle.
0:  If the entry cycle is less than $N_{t}$, reserve the VPs according to the feasible path, and assign the path to the robot; else, add the waiting time of the robot by $\tau_{c}$.
0:  Remove the robot from the candidate list and go to step 2.
0:  Record the newly generated VPs in the reservation table, labeled as unoccupied. Remove the VPs which have left the network. Output the assignment of feasible path.
So far, we have identified the essential components required to integrate heuristics into the RC-S framework: recording the status of VPs ever existed, and for each delivery task, finding a feasible path with an entry cycle, and reserving the VPs in the space-time map. To improve robot transporting efficiency and ensure the accessibility of each outlet, we allow up to three turns in each delivery route under this scheme. Compared with classic MAPF methods, it avoids potential path conflicts from the design stage and greatly reduces the search range while maximizing intersection capacity. The optimal strategy of RC-S requires selecting a combination of feasible path and entry time for each robot, and checking the occupancy status of VPs.
Algorithm 1 has a complexity of $O(N_{t}\cdot\lvert R\rvert)$ in each cycle. Moreover, when the algorithm is forcibly terminated midway, it can still produce a feasible solution, where robots that are not served will wait until the next cycle. The performance of Algorithm 1 will be validated in section 6.
5 Performance Analysis of RC-S
This section elaborates on the network capacity by analyzing the characteristics of RC-S and investigates how the system configurations affect system throughput. A throughput estimation formula is proposed, serving as a constraint for the planning model in section 7.
Two important metrics in traffic analysis (Daganzo,, 2010) are the total vehicular distance traveled per second of operation $\bar{m}(veh\cdot m/s)$, and the average travel distance $\bar{l}$. The relationship among system throughput $\tilde{T}_{O}$ and these two metrics is shown in equation (11):
$\tilde{T_{O}}=\frac{\bar{m}}{\bar{l}}$ (11)
Under the mechanism of RC-S, the first metric $\bar{m}$ is numerically consistent with the product of the number of occupied VPs, denoted by $n^{occupied}_{VP}$, and the speed of each VP, denoted by $v_{VP}$:
$\bar{m}=n^{occupied}_{VP}\cdot v_{VP}$ (12)
The speed of each VP could be obtained by expression (4)-(5) with the parameters of RC-S. For the number of occupied VPs, an intuitive idea is that it is related to the number of workers engaged in loading parcels in the loading zone and the spatial utilization under RC-S. We introduce two factors: workforce factor $\kappa$, representing the impact of the number of workers on the quantity of VPs that can be occupied; attenuation factor $\beta$, representing the utilization of VPs by the feasible paths generated by RC-S. In addition, the number of available robots $n_{r}$ also limits the upper bound of occupied VPs. Above all, the expression for calculating $n^{occupied}_{VP}$ can be formulated as:
$n^{occupied}_{VP}=\min\{\kappa\cdot\beta\cdot n_{VP},n_{r}\}$ (13)
where $n_{VP}$ denotes the total number of VPs in the network, which is constant given the scale of the network. According to the phase analysis in section 4.2, the RC-S scheme stipulates that the number of VPs is equal to the number of intersection nodes at any given time. Therefore, we obtain:
$n_{VP}=n_{h}\cdot n_{v}$ (14)
Now we investigate the forms of the two factors.
Workforce factor
This factor represents the impact of the number of workers on the quantity of VPs that can be occupied. For each aisle, the newly generated VPs at the entrance could be occupied by loaded robots when its corresponding loading station is active and worker is engaged in tasks. The quantity of such VPs is a critical determinant of the maximum throughput. Consider, for instance, a scenario where eight workers are stationed in an RSS with a $6+6$ aisle network. To ensure the balance of traffic flow and mitigate congestion at the aisle network’s periphery, workers are are evenly divided into four groups and positioned along the four edges of the site, as depicted in Figure 9. Four horizontal aisles and four vertical aisles are served by workers. We merge the area covered by these aisles and obtain the blue polygon.
Figure 9: The impact of number of workers on the VPs that can be occupied
For ease of discussion, the notation $\alpha$ will be employed to denote the ratio of workers to the network’s maximum worker capacity.
$\alpha=\frac{n_{w}}{n_{h}+n_{v}}$ (15)
The proportion of the blue region, noted as the workforce factor $\kappa$, could be obtained by:
$\kappa=1-{(1-\alpha)}^{2}$ (16)
We propose the following proposition regarding the relationship between the workforce factor and the number of occupied VPs:
PROPOSITION 1.
If either (i) the candidate feasible paths in RC-S include no more than two turns, or (ii) the average travel distance of a task is less than $(2-\alpha)min\{n_{v},n_{h}\}$, then an upper bound of the proportion of occupied VPs is $\kappa$;
The proof is presented in Online Appendix B. Proposition 1 provides with support for estimating the maximized network throughput. Next, we analyze the utilization rate of VPs using the RC-S scheme.
Attenuation factor
This factor represents the utilization of VPs by the feasible paths generated by RC-S. We begin with examine the path reservation strategy within the framework of RC-S . In a first-come-first-served basis, the system allocates to each newly requested sorting task an available robot, assigning it the currently shortest viable route. This robot then proceeds through the network along a predefined spatio-temporal path, ensuring a collision- and queue-free delivery within the aisle network. However, this spatio-temporal reservation mechanism obligates the robot to adhere strictly to its assigned spatio-temporal trajectory, precluding any alterations like deceleration or temporary halts. Such inflexibility can lead to underutilization of traffic capacity, as VPs that are momentarily unoccupied cannot always be sequentially linked to construct a continuous path if they do not intersect at the same point in time. To quantify the impact of this reduction, we introduce an attenuation factor $\beta$, representing the spatial utilization.
Figure 10: The ”pile up” process of spatiotemporal occupation under RC
Figure 11: The loss of capacity in different network
The attenuation factor $\beta$ rises with the increase of $(n_{h}+n_{v})$. Under the RC-S protocol, the aisle’s spatiotemporal occupation showcases a ”pile-up” effect, as depicted in Figure 10. The figure’s horizontal axis marks the spatial position within an aisle, while the vertical axis tracks discretized time intervals. Each robot’s transit through this path necessitates occupying a VP, visualized as a blue bar in the illustration. The collective occupation by different robots generates the ”pile-up” phenomenon. The capacity loss at any moment is represented by the unoccupied area beneath the upper boundary of this accumulation.
When comparing networks of varying scales, i.e., those with shorter versus longer aisle lengths, we maintain a continuous approach: assuming that new sorting tasks emerge with starting points evenly spaced along the boundary and target destinations uniformly spread across the network. In a stable operation, the flow distribution of paths under consistent control logic remains analogous across different network sizes, with variations only in total path lengths. Accordingly, the pattern of unoccupied spaces beneath the upper boundary remains consistent across scales. Figure 11 illustrates this by comparing capacity losses across aisles of diverse lengths. Longer aisles tend to exhibit extended periods of spatiotemporal occupation, potentially exacerbating capacity loss — the unoccupied intervals shown in Figure 11. In consequence, we can reasonably infer that the absolute amount of lost capacity increases quadratically with the number of aisles $(n_{h}+n_{v})$, and the proportion of the lost traffic capacity should grow linearly (because the number of available VPs increases linearly with the length of aisle). In summary, the expression of attenuation factor $\beta$ is as follows:
$\beta(n_{h},n_{v})=\frac{1}{a+b\cdot(n_{h}+n_{v})}$ (17)
where $a$ and $b$ are two parameters. Parameter $a$ is slightly greater than 1 and $b$ is relatively small, indicating that for a small-scale network, the attenuation factor should be close to 1 because there are fewer conflict points in one feasible path. We use simulation to obtain the real performance of RC-S and apply linear regression to calibrate the value of the parameters in the attenuation coefficient. Experiments show that a proper value is $a=1.4$, $b=0.012$. Note that, due to the presence of the continuity assumption, there is a certain bias in $\beta$ when the number of aisles is small. Specifically, in smaller networks, this factor may overestimate the system’s throughput, especially in scenarios where the activation rate of loading stations is low. This is because some path shapes may not exist in such networks. Experimental validation of this effect will be presented in the subsequent section. To this end, we derive the final expression of number of occupied VPs:
$n^{occupied}_{VP}(\alpha,n_{r})=\min\{\frac{1-{(1-\alpha)}^{2}}{a+b(n_{h}+n_{v})}\cdot n_{h}n_{v},\ n_{r}\}$ (18)
LEMMA 1.
$n^{occupied}_{VP}(\alpha,n_{r})$ is non-decreasing and concave for all $\alpha$ and $n_{r}$ in its domain.
The proof of Lemma 1 can be accomplished by verifying that the partial derivatives of the two terms in the minimum function are non-negative and the Hessian matrices are both semi-negative definite. Lemma 1 indicates that opening more loading stations will increase the capacity of VPs within the network, thereby raising the upper limit of total throughput. However, the marginal returns of constructing new loading stations will gradually diminish.
Average travel distance
We now derive the average travel distance $\bar{l}$ of a sorting task, as the remaining part of the throughput estimation formula. Based on the mechanism of RC-S, an appropriate selection of paths is based on the principles of evenly distributed sorting demand, workload balancing of workers, and minimizing travel distance. Workload balancing principle ensures that loading stations do not experience starvation, thus preserving the stability of the system. It also aligns better with real-world scenarios. To maintain a balanced flow of robots between loading stations, we arrange the workers at the centers of each side, covering a length equal to $\alpha$ times the side length. Figure 12 illustrates this layout, where outlets are divided into four types of areas based on their positions. Sorting demand are uniformly and continuously distributed across the entire area. As mentioned earlier, each time a robot turns, it occupies additional space in the VP fleet, so it is necessary to restrict the maximum permissible robot’s turning times in one delivery. However, from the diagram, it can be observed that to ensure the outlets in area 3 are reachable from each loading station, paths with at least three turns are required. According to our simulation experiments, introducing paths with four or more turns did not effectively improve the sorting throughput. Therefore, we prioritize paths with fewer turns and do not consider paths that require more than three turns. We denote the path length serving area $i$ as $l_{i}$, with the proportion of this area represented by $p_{i}$. We will next present the rationale for their calculation and the formulas. For a detailed derivation process, please refer to the appendix B.
Figure 12: Dividing areas according to the layout
l
Figure 13: Service paths for different categories
Firstly, for area 1, servicing the outlets requires one change of direction, as shown in Figure 13(a1). Based on the balance of workload, the turning points should have an equal probability of occurring within the central area. We then perform a weighted accumulation based on the length and then take the average. Consequently, we obtain expressions for the area proportion and the expected path length:
$p_{1}$ $=\alpha(1-\alpha)$ (19)
$E[l_{1}]$ $=2D\cdot[\frac{\frac{9+a^{2}}{6}(n_{h}^{2}+n_{v}^{2})-n_{h}n_{v}-\frac{1}{3}}{n_{h}+n_{v}}+1]$ (20)
For paths that pass by the areas of area 2, their endpoints are located at either the lower or upper loading stations, as shown in Figure 13(b1) and (b2), respectively. Let $(x_{1},y_{1})$, $(x_{2},y_{2})$ denote the coordinates of the two turning points in a path. $x_{1}$ and $x_{2}$ are uniformly distributed over the length covered by the workers, while the $y_{1}$ and $y_{2}$ are identical and uniformly distributed across the area. Notably, when x1 = x2, the path actually does not involve any turns. Let the robot choose the nearest top or bottom exit point when making the second turn. The expressions for $p_{2}$ and $E[l_{2}]$ are as follows:
$p_{2}$ $=\alpha(1-\alpha)$ (21)
$E[l_{2}]$ $=2D\cdot[\frac{\alpha(n_{h}^{2}+n_{v}^{2})}{3(n_{h}+n_{v})}-\frac{2}{3\alpha(n_{h}+n_{v})}+\frac{3n_{h}n_{v}}{2(n_{h}+n_{v})}]$ (22)
For Area 3, as previously mentioned, only paths involving three turns can service it. Area 3 includes four individual small areas. The two small areas on the bottom, as their service paths shown in Figure 13(c1), have a length equal to the paths with one turn in Figure 13(a1); the two areas on the top, with service paths depicted in Figure 13(c2), involve two extra segments of path in the y-direction (or in the x-direction when the starting point is on the left or right side). As a result, the expressions for $p_{3}$ and $E[l_{3}]$ are as follows:
$p_{3}$ $={(1-\alpha)}^{2}$ (23)
$E[l_{3}]$ $=2D\cdot(\frac{n_{h}+n_{v}}{2}+\frac{(1+\alpha)}{4}\cdot\frac{n_{h}n_{v}}{n_{h}+n_{v}})$ (24)
Finally, for Area 4, which is in the center, we assume it is equally served by the paths of area 1 and area 2.
$p_{4}$ $={\alpha}^{2}$ (25)
$E[l_{4}]$ $=\frac{1}{2}E[l_{1}]+\frac{1}{2}E[l_{2}]$ (26)
In summary, the average travel distance of a sorting task can be obtained by the following weighted summation:
$\bar{l}(n_{h},n_{v},\alpha)=\sum_{i=1}^{4}p_{i}E[l_{i}]$ (27)
Now, we further investigate the properties of equation (27).
LEMMA 2.
For uniformly distributed sorting demands, if the length-to-width ratio is less than 2 and $min\{n_{h},n_{v}\}\geq 4$, (i) as $\alpha$ increases, $\bar{l}(n_{h},n_{v},\alpha)$ decreases initially and then increases; (ii) when the site size is fixed ($n_{h}\cdot n_{v}=\textit{constant}$), a square-shaped site has a smaller average travel distance; (iii) when the site is square-shaped, the range of average travel distance could be obtained by $2D\cdot n_{h}<\bar{l}(n_{h},n_{v},\alpha)<2D\cdot\frac{9}{8}n_{h}$.
Proofs for Lemma 2 is provided in Appendix B. The condition within the lemma indicates that the conclusion is applicable to an RSS where the length-to-width ratio does not exceed two. This is a reasonable setup in practice, as excessively elongated sites accommodate fewer outlets under the same area. Lemma 2(i) states that the efficiency of sorting by robots exhibits a trend of first increasing and then decreasing with the increase in the number of workstations. This is because, under the worker load balancing criterion, the average distance from the corner loading stations to each outlet is greater than the stations located in the middle of the four sides. Lemma 2(ii) suggests that, under the permissible conditions, designing the aisle network in a square shape reduce the demand for robots. As aforementioned in section 3, the density of outlets remains constant, thus the distribution of outlets in RSS does not need to be altered as well. Lastly, Lemma 2(iii) provides the range of average distance to be covered in a single delivery within RC-S. This will serve as a foundation for the efficiency analysis in the following sections.
Combining equations (15)-(27), the estimation formula of sorting throughput under given configuration $(n_{h},n_{v},n_{w},n_{r})$ is as follows:
$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})=\frac{D\cdot n^{occupied}_{VP}(\frac{n_{w}}{n_{h}+n_{v}},n_{r})}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{w}}{n_{h}+n_{v}})}$ (28)
If we fix the values of $n_{h}$ and $n_{v}$, the upper bound of system throughput could be calculated by setting the value of $n_{r}$ to infinity and assigning workers to all loading stations:
$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})\leq\frac{D\cdot n^{occupied}_{VP}(\frac{n_{l}}{n_{h}+n_{v}},\infty)}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$ (29)
We use $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ to denote this upper bound in the right side of inequality (29). Here we investigate the properties of this bound.
PROPOSITION 2.
$\frac{\partial\tilde{T}_{M}(\cdot)}{\partial n_{l}}$ has one zero point within the valid range of $n_{l}$. Denote the value as $n_{l}^{(0)}$, then $\tilde{T}_{M}(\cdot)$ is monotonically increasing in the interval $(0,n_{l}^{(0)}]$ and monotonically decreasing in the interval $(n_{l}^{(0)},n_{h}+n_{v}]$.
The proof of Proposition 2 can be derived by incorporating Lemma 1 and Lemma 2, which are depicted in detail in Appendix B. Proposition 2 reveals that under the RC-S robot control scheme, the maximum throughput has a critical point, and this critical point is not achieved when all loading stations are activated, due to workload balancing and uniformly distributed demands. When the proportion of active loading stations approaches 1, the growth rate of $n^{occupied}_{VP}$ will be slower than the growth rate of the service path length, resulting in a decrease in throughput. We will verify the formula through experiments in the next section.
6 Numerical Validation
6.1 Comparative analysis: RC-S versus benchmark traffic management framework
In this section, we compare our proposed RC-S against cooperative A* (CA*) (Silver,, 2021) for traffic management in a warehouse, to better illustrate the superiority of RC-S in terms of AGV service distance and system throughput. CA* is a framework based on a simple prioritized-planning scheme: each agent is first assigned a unique priority, and based on the order of priorities, algorithm find the shortest path for each agent that avoids conflicts with agents of higher priority. We use SIPP (Phillips and Likhachev,, 2011), an efficient variant of location-time A*, as the lower-level solver of CA*. Moreover,other well-known MAPF algorithm frameworks, such as PBS (Ma et al.,, 2019) and RHCR (Li et al.,, 2021), are found less efficient compared to CA* in large-scale settings ($n_{r}>100$) of our experiments, primarily due to the increased likelihood of cycle conflicts within the RSS framework, and thus are excluded from this comparative analysis.
Table 2: RC-S parameters in simulation
$D(m)$ $\tau_{e}(s)$ $\tau_{c}(s)$ $v_{VP}(m/s)$
1 0.5 2 2
We compare the performance in two scenarios, 12+12 aisles and 20+20 aisles. To ensure uniformity in the experimental setup, we use directed maps in all tests to prevent swapping conflicts. We activate all possible loading station locations for the experiments, and parcel targets are evenly distributed among outlets. The RC-S parameters, detailed in Table 2, are based on sorting robot data from the Geekplus company. For a fair comparison of algorithm performance, robots move at a constant speed of 2m/s, matching the speed of VPs in RC-S. The simulations do not consider acceleration and deceleration phases. We conduct 50 repeated experiments for each scenario. In each experiment, we set a warm-up period of 30 minutes and then record the system performance for 60 minutes of continuous operations. Our algorithms and simulations are coded in Python 3.11, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM. All processes are run single-threaded.
Table 3: Results on 12+12 aisle network
$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time per cycle(ms)
RC-S CA* RC-S CA* RC-S CA*
40 8.83 4.72 11.73 14.17 5.67 5.99
80 15.12 9.03 12.92 14.57 6.80 16.85
120 17.45 12.40 13.89 15.24 8.50 35.37
160 18.06 14.32 14.26 15.79 9.31 55.31
200 18.33 14.94 14.51 16.03 9.63 63.36
Table 4: Results on 20+20 aisle network
$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time per cycle (ms)
RC-S CA* RC-S CA* RC-S CA*
50 7.68 3.94 18.85 21.75 15.53 11.61
100 14.91 7.79 19.33 21.95 15.79 29.94
200 26.82 15.07 20.77 22.54 19.76 99.61
300 32.32 22.58 21.48 23.33 29.27 226.76
400 33.66 24.37 23.41 23.92 37.40 363.45
Tables 3 and 4 report the system throughput, average service time and average run-time. The comparative results indicate that the run-time of RC-S is consistently lower than that of CA*, and in larger-scale cases, it is even less than $1/10$ of the latter. Moreover, RC-S always assigns shorter service paths to robots, resulting in an average service time reduction of 10.3% compared to CA*. In most experiments, RC-S demonstrates superior throughput compared to CA*, with the exception observed in a 20+20 aisles scenario with $n_{r}=400$. This observation underscores the suitability of the RC-S algorithm for environments like RSS, characterized by a compact network and numerous outlets as obstacles. The difference in algorithm performance is primarily due to the following improvements: (1) RC-S optimizes by selecting from candidate spatio-temporal paths, while CA* and similar cell-based search algorithms consider variable waiting times at each cell and allow robots to return to previously visited cells; (2) RC-S limits the total number and maximum density of robots and is free of conflicts, while high robot density around outlets in benchmark algorithm leads to significant resource expenditure on conflict resolution.
6.2 Throughput estimation formula validation
This section validates the throughput estimation formula proposed in section 5 across different active ratios of loading stations. We consider 5 scenarios with different network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set as the number of activated VPs in the aisle network, plus 5 per loading station to ensure the stability of operation. Other experimental parameters are specified in section 6.1 in detail.
Figure 14: Analytical formula compared with simulation results.
Figure 15: Robot flow distribution in simulation, $n_{h}=n_{v}=20$
Figure 14 shows the trend of the average service distance as a function of the number of active workstations when $n_{h}=n_{v}=20$. It indicates that the formula and simulation results present the similar trend of first decreasing and then increasing. This suggests that the service frequency of each robot is highest when about half of the loading stations are activated. Figure 14 shows the relative error between formula and simulation results under five scenarios. It can be seen that the error is always within ±$\pm 8\%$, indicating a good fit.
Figure 14 shows the trend of the system throughput as a function of the number of activated workstations when $n_{h}=n_{v}=20$. The trend demonstrates a clear marginal effect of the number of activated loading stations: when the system is operating at the maximum capacity, there is a certain degree of resource waste. Figure 14 shows the relative error between formula and simulation results for five cases, which remains within ±$\pm 15\%$. Figure 15 depicts the average traffic flow distribution during a stable operation of the system in simulation with the maximum number of workers. The darker the color of the aisle segments, the higher the traffic load. We notice that the corner loading stations have almost no incoming or outgoing traffic flow. This indicates that Algorithm 1, in practical execution, cannot effectively balance the distribution of robots in the loading stations in the extreme situation.
6.3 Efficiency underestimating in queueing models
Queueing models have been widely used in analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different work pieces (John A. Buzacott,, 1993). In robotic systems, robots function similarly to pallets in traditional manufacturing systems, transferring work pieces from one machine to another for processing. However, we find that queueing models could not capture the efficiency of robots accurately when the number of robots has not reached the network’s capacity limit. We apply the CQN model established in Zou et al., (2021), where a two-tier RSS model with closest Robot-to-Loading-Station assignment rule is modified to meet the assumptions in section 3. Eight scenarios of large-scale network with few robots are tested, using the same experimental setup as in section 6.2. The results are shown in Table 5, where Eq.(28) represents the proposed estimation formula. The CQN model significantly underestimates system throughput in scenarios with few robots. It lacks the capability to accurately depict traffic capacity under low flow conditions with RC-S control and overestimates the impact of conflicts under low-demand conditions, thereby underestimating individual robot efficiency. Although the CQN model provides accurate estimates when the network approaches capacity, it results in unnecessary expenses for surplus robots during off-peak periods.
Table 5: Error in throughput estimation under RC-S control
$n_{h}+n_{v}$ $20+20$ $24+24$
$n_{r}$ 50 100 200 300 50 100 200 300
Throughput (×$\times 10^{3}$/h) Simulation 8.43 16.25 28.60 32.90 7.27 14.30 27.01 36.49
CQN 5.98 11.87 22.98 31.47 4.96 9.88 19.52 28.61
Eq.(28) 7.83 15.66 31.32 33.86 6.55 13.10 26.19 39.08
Relative error (%) CQN 29.03 26.95 19.65 4.34 31.82 30.92 27.73 21.61
Eq.(28) 7.12 3.65 -9.52 -2.92 9.92 8.41 3.04 -7.10
we also conduct a experiments on a real case of Deppon Express described in Zou et al., (2021), where the network scale is $n_{h}+n_{v}=18+6$, with 108 outlets, 6 loading stations and 170 robots. The The aisle width $D$, maximum robot speed $v_{max}$, and worker loading rate $r_{l}$ were set according to the data provided by Deppon Express. We compared the performance of RC-S with the results from Zou et al., (2021), as shown in Table 6, where $T_{D}$, $T_{CQN}$ and $T_{RCS}$ represent the throughput of the Deppon system, the CQN model and the simulation under RC-S scheme, respectively. Combined with the experimental results in section 6.1, RC-S scheme significantly improves the operational efficiency of robots, thereby enhancing the overall system throughput. Meanwhile, the CQN model effectively captures the congestion in the RSS system implemented by Deppon Express, but underestimates the traffic capacity of the network when the RC-S scheme is applied.
Table 6: Real Case Validation
$D(m)$ $r_{w}(/s)$ $v_{max}(m/s)$ $\tau_{c}(s)$ $v_{VP}(m/s)$ $T_{D}(/h)$ $T_{CQN}(/h)$ $T_{RCS}(/h)$
0.6 0.42 2 1.2 2 7,163 6,907 8,758
6.4 Evaluating RC-S across diverse scenarios
In real sorting centers, due to cost constraints and demand arrival rates, loading stations are often not spread across the entire perimeter of the RSS, resulting in only partial availability of network entrances and exits. The placement of outlets is also restricted by site factors, such as load-bearing columns or the presence of certain equipment. Additionally, demand arrival rates are heterogeneous and are not uniformly distributed across each outlet. In this section, we evaluate the performance of RC-S across various application scenarios and test the impact of different system configurations on efficiency by simulation. For each set of experiments, we evaluate the overall performance across 50 test runs. Each run includes a 30-minute warm-up period, followed by an observation period where average data is recorded over one hour.
We first compared the impact of different numbers of active loading stations on average throughput. We consider five scenarios with varying network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set to be sufficient in each scenario, ensuring that it does not become a bottleneck for sorting process. Other RC-S parameters are specified in detail in Table 2. Figure 16 shows the trend of throughput across the five scenarios. It can be observed that large-scale networks not only have a higher throughput upper bound, but their slope—representing the average efficiency per loading station—is also generally higher. This is because, under RC-S, larger networks provide more space to distribute incoming traffic across different aisles, thereby reducing congestion at each entrance.
Figure 16: The effect of different numbers of active loading stations
(a) Division of outlet areas in the network
(b) RC-S performance in different scenarios
Figure 17: The effect of closing outlets in the network
We next evaluate the effect of closing some outlets on system throughput. The network size is set to $n_{h}=n_{v}=20$, with other experimental parameters remaining constant. We consider four scenarios for closing outlets: (1) randomly closing outlets; (2) closing outlets in the corners; (3) closing outlets along the edges; (4) closing outlets in the center. To control variates, the number of closed outlets in each scenario is set to 1/9 of the total. Scenarios (2), (3), and (4) correspond to areas 1, 2, and 3 in Figure 17(a), respectively. Results from 50 repeated experiments for each scenario are shown in Figure 17(b). The green dashed line represents the average throughput when all outlets are open. It can be observed that randomly closing outlets has no significant impact on system throughput. Closing outlets in areas 1 and 2 results in a decrease in average throughput, while closing outlets in area 3 leads to a slight increase in average throughput. It is intuitive that closing outlets in a specific area causes an uneven distribution of demand, leading to imbalanced network traffic and increased traffic load in other areas of the network. However, from another perspective, it reduces robot turns within the closed area, thereby improving the utilization of spatio-temporal resources. Since most shortest paths pass through the central area, closing central outlets in scenario 4 results in higher throughput compared to the normal case.
Figure 18: The effect of heterogeneous demand distribution
In addition, we evaluate the effect of heterogeneous demand distribution on system efficiency. In real sorting centers, sorting destinations with higher demand arrival rates are often assigned multiple outlets to minimize differences in the visit frequency for each outlet; however, this cannot entirely eliminate demand heterogeneity. We assume there are two types of outlets in the RSS: one type has an average demand arrival rate twice that of the other, with both following a normal distribution. For outlets with extremely low arrival rates, the situation is similar to the previous experiments involving closed outlets, and their impact can be minimized by random assignment. The two types of outlets are spatially uniformly distributed throughout the network. Keeping other experimental parameters constant, we adjust the proportion of the two types of outlets, and the system performance are shown in Figure 18. The results indicate that with a fixed total demand and varying proportions of high-demand outlets, RC-S maintains efficiency, and the impact of demand heterogeneity on system throughput is negligible. This is achieved by leveraging demand arrival rates for different destinations and balancing the visit frequency to each area, ensuring a well-distributed traffic flow across the network.
Table 7: RC-S performance in large-scale system
$n_{h},n_{v}$ Throughput Avg. path Avg. time Max. time # Feasible paths
($\cdot{10}^{4}$/h) length (m) per cycle (s) per cycle (s)
22 3.871 27.832 0.0359 0.0942 118,136
24 4.331 30.062 0.0480 0.1255 167,084
26 4.699 32.517 0.0614 0.2393 229,888
28 5.156 34.734 0.0836 0.2436 308,948
30 5.517 37.261 0.1004 0.2820 406,856
[…]
7 Layout Design Optimization
The layout design problem aims at seeking for a balance between facility costs and long-term operating costs. A larger facility will achieve higher maximum throughput; however, it will also increase construction and equipment costs. Considering transportation costs and supply stability, we analyze the following scenario: The warehouse signs a long-term RaaS contract with a robotics company, with each year divided into several operating periods denoted by $\sigma$. The proportion of each period’s duration within the year is represented by $\theta^{\sigma}$, and the planned throughput level during each period is $T^{\sigma}$. The facility costs $C_{f}$ and operations costs $C_{o}$ are calculated as follows.
$C_{f}=P_{s}\cdot[2D\cdot(n_{h}-1)+W_{w}+W_{l}]\cdot[2D\cdot(n_{v}-1)+W_{w}+W_{l}]+P_{l}\cdot n_{l}$ (30)
$C_{o}=\sum_{\sigma\in\mathcal{S}}\theta^{\sigma}(P_{w}n_{w}^{\sigma}+P_{r}n_{r}^{\sigma})$ (31)
Parameter $P_{s}$ and $P_{l}$ in equation (30) denote the discounted site rental cost per square meter and equipment cost per loading station, while $P_{w}$ and $P_{r}$ in equation (31) denote the discounted unit labor cost and unit rental cost of robot. The objective of the problem is modeled as follows:
$C_{d}(T^{\sigma},\theta^{\sigma},N_{o})=C_{f}+C_{o}$ (32)
where parameter $N_{o}$ denotes the minimum number of outlets that should be covered. The trade-off between these two types of costs in the objective function is the focus of the site planning stage. We formulate the layout design problem (LDP) as following integer programming model.
$\underset{n_{h},n_{v},n_{l},n_{w}^{\sigma},n_{r}^{\sigma}}{\min}\ C_{d}$
[…]
$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{\sigma},n_{r}^{\sigma})\geq T^{\sigma}\qquad\forall\sigma\in\mathcal{S}$ (33)
$(n_{h}-1)(n_{v}-1)\geq N_{o}$ (34)
$n_{l}\leq n_{h}+n_{v}$ (35)
$n_{w}^{\sigma}\leq n_{l}\qquad\forall\sigma\in\mathcal{S}$ (36)
$n_{h}=2k_{h}\ ,\quad k_{h}\in\mathcal{Z}_{+}$ (37)
$n_{v}=2k_{v}\ ,\quad k_{v}\in\mathcal{Z}_{+}$ (38)
$n_{w}^{\sigma},n_{r}^{\sigma}\in\mathcal{Z}_{+}^{2\left|\mathcal{S}\right|}$ (39)
Constraint (33) ensures that the system throughput meets predicted sorting requirements in different periods. Constraint (34) restricts the lower bound of the site size to accommodate an adequate number of outlets. Constraint (35) represents the maximum number of loading stations accommodated in the aisle network under RC-S. Constraint (36) limits the maximum number of workers to be stationed. Constraints (37)-(38) ensure the number of aisles even. The remaining constraint (39) requires the variables to be positive integers.
It is difficult to solve LDP because of the non-linearity of the inequality constraints, especially for large-scale problems. In this research, we apply the method of penalty successive linear programming (PSLP) to solve the LP-relaxation of LDP, which exhibits good robustness and convergence properties for large-scale problems (Bazaraa et al.,, 2013). Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem along with the penalty function, and utilizes the concept of trust region (updated at each iteration) to control the step size. To ensure the constraints continuous and differentiable, we introduce additional constraints requiring that the number of robots is less than the number of available VPs, shown as follows.
$n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\quad\forall\sigma\in\mathcal{S}$ (40)
At the end of the algorithm, the optimal solution of the relaxed problem is converted into a feasible solution for the original problem by rounding up. The details of PSLP algorithm are shown in appendix C.
8 Numerical Examples for Layout Design
In this section, we conduct a sensitivity analysis on the site rental cost per square meter $P_{s}$, and labor cost per man-month $P_{w}$, focusing on the trend of total costs under different scenarios as the sorting throughput level changes. Due to increased competition in the robot market, prices remain relatively stable across different companies. Consequently, we do not investigate variations in robot rental cost during the experiments, nor do we examine equipment cost at loading stations. We consider a 5-year investment plan with a monthly interest rate $\gamma_{0}=0.5\%$, and then the discounted costs of different components in LDP are calculated by:
$P_{i}=\sum_{t=0}^{60}\frac{M_{i}}{{(1+\gamma_{0})}^{t}}\ ,\quad i=s,l,w,r$ (41)
where $M_{i}$ denotes the unit cost of component $i$ in a month. We estimate the monthly warehousing rental cost using JD Logistics’ financial reports for the first three quarters of 2023. Labor cost is sourced from data related to warehousing job recruitment in Beijing, obtained from website for job posts (58.com). The data regarding the equipment cost of loading stations and the rental cost of robots are obtained from Geekplus company.
We assume that the warehouse only expands its sorting capacity during a few major shopping events, meaning there are two typical levels of sorting demand, namely the average off-peak season demand ${T}^{L}$ and the average peak season demand ${T}^{H}$, with ${T}^{L}=0.8\cdot{T}^{H}$. The ratios of the two periods are set as $(\theta^{L},\theta^{H})=(5/6,1/6)$, which means that each quarter typically includes a two-week-long shopping event. The default values of parameters is shown in table 8 and the configurations of RC-S keeps the same with table 2. Our algorithms are coded in Python 3.11 with the solver Gurobi Optimizer 9.0, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM. All processes are run single-threaded.
Table 8: Default values of parameters in section 8
$M_{s}(CNY/m^{2}\cdot mo)$ $M_{l}(CNY/mo)$ $M_{w}(CNY/mo)$ $M_{r}(CNY/mo)$ $W_{w}(m)$ $W_{l}(m)$
10 400 5,000 200 5 5
8.1 Sensitivity Analysis on the site rental cost
We begin by investigating the impact of site rental cost under two different scenarios: one with $N_{o}=100$ (scenario 1) and another with $N_{o}=400$ (scenario 2). The monthly rental cost of site floor space per square meter varies from $10\ CNY$ to $30\ CNY$. The results are presented in Figure 19(a) and Figure 19(b). The findings reveal that:
(a) $N_{o}=100$ (scenario 1)
(b) $N_{o}=400$ (scenario 2)
Figure 19: Comparison of total costs and facility costs proportion on different site rental cost
Table 9: Results of scenario $M_{s}=10,N_{o}=100$
${T}^{H}$ $(n_{h},n_{v})$ $(n_{w}^{L},n_{r}^{L})$ $(n_{w}^{H},n_{r}^{H})$ $C_{f}$ $C_{o}$ $C_{d}$ Proportion of
(×$(\times{10}^{6})$ (×$(\times{10}^{6})$ (×$(\times{10}^{6})$ site rental cost (%)
3000 (10,12) (2,10) (3,13) 0.53 0.67 1.20 38.80
6000 (10,12) (4,21) (5,26) 0.57 1.31 1.88 24.78
9000 (10,12) (7,33) (9,42) 0.65 2.26 2.92 15.96
12000 (12,14) (9,51) (13,67) 0.87 3.07 3.94 15.20
15000 (14,16) (12,75) (18,98) 1.12 4.20 5.32 14.07
18000 (18,18) (15,109) (23,142) 1.48 5.44 6.92 14.54
21000 (20,22) (18,147) (28,194) 1.88 6.72 8.60 15.09
24000 (24,26) (21,199) (32,261) 2.41 8.11 10.52 16.60
27000 (26,26) (25,236) (44,315) 2.79 9.91 12.70 14.74
30000 (30,30) (29,302) (49,401) 3.42 11.72 15.14 15.88
  • When the target throughput level is low, the total costs in scenario 1 is significantly lower than that in scenario 2. This is because, to accommodate more outlets, the latter requires renting a larger initial space. In both scenarios, the curve of the optimal system cost exhibits a turning point $T^{*}$. When $T^{H}>T^{*}$, the cost escalation rate increase significantly. The turning point indicates that the initial space in the current scenario is no longer sufficient to accommodate a larger throughput, necessitating an expansion of the aisle network scale. In scenario 1, this turning point occurs at $T^{*}=9,000$; whereas in scenario 2, the turning point is at $T^{*}=21,000$.
  • The proportion of facility costs increases as $M_{s}$ rises, particularly remarkable at lower sorting throughput levels. As the throughput level becomes higher, this proportion gradually decreases and stabilizes after the turning point $T^{*}$ in each case. By observing the system configurations in each dataset, we find that this phenomenon arises because, at higher throughput demands, efficiency improvements necessitate expanding the space, which leads to the increase in the site rental cost. Table 9 shows the optimal system configuration and corresponding costs in the case of $M_{s}=10$ and $N_{o}=100$. When $T^{H}>T^{*}$, the optimal solutions show not only assigning more workers and robots to loading stations but also simultaneously expanding the size of the network to alleviate traffic pressure.
  • Comparing the data for $T^{H}=15,000$ and $T^{H}=30,000$ in Table 9, the total costs for the latter are nearly three times those for the former. This indicates that due to the simultaneous growth of $C_{f}$ and $C_{o}$, the total costs growth rate of the RSS system exceeds the throughput growth rate.
From the results of study in site rental cost, the derived managerial insights are summarized as follows.
INSIGHT 1.
The advantage of RSS is evident in its low initial investment in scenarios with both throughput demand and sortation category low.
INSIGHT 2.
High density lead to decreased efficiency of workers. Instead of running RSS at full capacity, it is more cost-effective to appropriately expand the site size and reduce the proportion of activated loading stations.
(a) $N_{o}=100$ (scenario 1)
(b) $N_{o}=400$ (scenario 2)
Figure 20: Comparison of total costs and labor costs proportion on different unit labor cost
8.2 Sensitivity Analysis on the labor cost
In a similar manner as our previous experiments, we now explore the influence of different labor costs on the total costs in two scenarios with $N_{o}=100$ and $N_{o}=400$. The monthly labor cost of each worker varies from $5,000\ CNY$ to $12,000\ CNY$. The proportions of labor costs and robot rental costs are both visualized in Figure 20(a) and Figure 20(b). Our observations are concluded as follows:
  • As $T^{H}$ increases, the proportion of labor costs gradually rises and eventually stabilizes. In every scenario, the proportion of labor costs in the optimal solution for $T^{H}=30,000$ exceeds 50%.
  • The rental cost of robots gradually increases with the level of throughput in each scenario. This is because, to maintain traffic efficiency, the number of robots needs to increase with the expansion of the network. Meanwhile, as discussed in the performance analysis in section 5, the rate of traffic capacity loss also increases, resulting in a considerable robot rental cost.
  • Comparing the two scenarios, it is evident that in larger-scale sorting sites (scenario 2), the initial increase in the proportion of robot rental cost is faster, making it challenging for RSS systems to balance the advantages of high scalability and low cost.
From the result of study in labor cost, we summarize the following managerial insight:
INSIGHT 3.
At lower levels of throughput, facility costs are the primary expense of RSS, whereas at higher levels of throughput, labor costs becomes the primary expense.
Additionally, the cost structure undergoes notably changes with varying sorting demands. In scenarios with high sorting demand, the proportion and growth rate of robot rental cost become more significant. This reveals a distinctive feature of RSS systems: Robots, as carriers of parcels, become less efficient in larger sites because they have to cover increased distances. Merely enhancing management methods is insufficient to tackle the challenge of diminishing traffic efficiency. Today’s commonly used conveyor sorters, equipped with high-power motors, necessitate substantial initial investments but experience only slight increments in total costs as sorting demands grow (Zou et al.,, 2021). Furthermore, the optimal system configurations of a conveyor sorter are not easily affected by fluctuations in the unit price of cost components (Russell and Meller,, 2003), which makes it suitable for long-term operation in large-scale sorting scenarios. Therefore, we conclude the following insight regarding the application of RSS:
INSIGHT 4.
Due to the traffic issues of robots, RSS is suitable for application in small-scale scenarios, such as distribution centers at the end of the supply chain.
9 Conclusion
This study conducts a comprehensive analysis of the RSS in modern warehouses, encompassing the underlying robot traffic management, efficiency analysis, as well as the cost composition and layout design of the system. By incorporating the throughput estimation, this research sheds light on the prediction of performance and resource allocation in modern warehouses or factories that utilize numerous robots, promoting the ongoing development of automation in logistics industry.
In the operations stage, efficient robot scheduling emerges as a crucial element for unlocking the business value of RSS. We propose an innovative RC-S scheme, which serves as a framework for managing a considerable number of robots simultaneously and efficiently. We provide a detailed description of the composition of phases and the method for setting parameters in RC-S. Additionally, we present a mathematical programming model FPA to minimize the travel costs at each cycle, and a heuristic algorithm as online solver. Subsequently, a theoretical analysis of the efficiency of RC-S is conducted, exploring the impact of various system configurations on throughput. In the validation section, we first compare our proposed traffic management framework with classical cooperative A* algorithm. The simulation results indicate that our control method achieves a higher level of performance and computational efficiency. Furthermore, we validate the throughput estimation formula. Results show that queueing models can experience significant distortion in some scenarios, while our model fits the real traffic flows of robots more precisely.
In the site planning stage, warehouse managers need to make decisions that involve a trade-off between the initial investment of facility $C_{f}$, and the discounted operations costs $C_{o}$. We analyze the cost composition of the RSS, and propose a layout design optimization model LDP that minimizes the total system costs. Through a detailed sensitivity analysis examining the cost dynamics of site rental and labor, we investigate cost proportions under distinct throughput levels. Specifically, at lower throughput, the rental cost for the site represents a considerable proportion of the total costs. However, at higher throughput, the expenditure on labor emerges as the predominant cost element. Results also validate the discussions on the properties of our model. The key insights can guide managers in understanding the investment and returns of applying the RSS, thereby reducing the total costs of the warehouse.
For future work, there are two directions to explore. Firstly, optimizing layout design considering non-uniform demand and various outlet distributions would further improve the model’s effectiveness. Secondly, extending the analysis to cover the entire process from order fulfillment to sorting in modern warehouses would provide valuable insights for enhancing system performance.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to check for grammatical errors and improve readability. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Appendix A: Main notations
Table A-1: Summary of main abbreviations
Abbreviations Explanation
RSS robotic sorting system
RMFS robotic mobile fulfill system
RaaS Robot-as-a-Service
MAPF multi-agent path finding problem
SOQN semi-open queueing network
OQN open queueing network
RC-S rhythmic control for sorting scenario
VP virtual platoon
FPA feasible path assignment problem
LDP layout design problem
PSLP penalty successive linear programming
Table A-2: Notations and explanation
Notations Explanation
Sets
$\mathcal{G}$ graph constituted of all nodes and links
$\mathcal{V}_{c}$ set of conflict nodes
$\mathcal{V}_{u}$ set of unloading nodes
$\mathcal{V}_{e}$ set of entrance/exit nodes
$\mathcal{E}$ set of all edges
$\mathcal{L}$ set of all loading stations
$\mathcal{O}$ set of all outlets
$\mathcal{C}$ set of all cycles
$\mathcal{R}$ set of all feasible path
$\mathcal{R}(i,j,k)$ set of feasible path connecting loading station $i$ and $j$, passing outlet $k$
$\mathcal{R}_{i},k$ set of feasible path starting at loading station $i$ and passing by outlet $k$
^$\hat{\mathcal{R}}_{i}$ set of feasible path ending at loading station $i$
$\mathcal{I}_{x}$ set of locations of loading stations on the top and bottom sides
$\mathcal{I}_{y}$ set of locations of loading stations on the left and right sides
$\mathcal{S}$ set of operating periods
Parameters
$D$ length of each grid in the sorting zone
$W_{w}$ width of waiting zone
$W_{l}$ width of loading zone
$C_{f}$ the facility costs of RSS
$C_{o}$ the operations cost of RSS
$C_{d}$ the total costs of RSS
$P_{f}$ cost of facility per square floor space
$P_{l}$ cost of equipment in one loading station
$P_{w}$ cost of labor per person per month
$P_{r}$ cost of robot per vehicle per month
${T}^{\sigma}$ target throughput level in operating period $\sigma$
$\theta^{\sigma}$ Ratio of operating period $\sigma$
$N_{o}$ minimum number of outlets required to meet the sorting category demands
$\gamma$ weight of operations costs in optimization model
$\tau_{c}$ time interval of VPs in RC-S
$\tau_{e}$ travel time of VPs on each link in RC-S
$v_{VP}$ speed of VPs
$v_{max}$ maximum speed of robots
$c_{max}$ maximum acceleration/deceleration of robots
$r_{l}$ maximum loading rate of a loading station
$c_{i,k}^{r}$ travel time of feasible path $r\in R_{i,k}$
$\hat{c}_{i,k}$ penalty of delay of a sorting task from loading station $i$ to outlet $k$
$\delta_{i,k}^{r,\nu,l}$ incidence between feasible path $r\in R_{i,k}$, node $\nu$ and cycle $l$
$N_{\nu}^{l}$ remaining capacity in node $\nu$ in cycle $l$
$d_{i,k}$ sorting demand from loading station $i$ to outlet $k$
Variables
$x_{i,k}^{r}$ decision on whether feasible path $r\in R_{i,k}$ is reserved in current cycle
$\hat{x}_{i,k}$ decision on whether the robot located at loading station $i$ should wait until the next cycle, with the target outlet $k$
$n_{h},n_{v}$ number of horizontal and vertical aisles in sorting zone
$n_{l}$ number of loading stations in loading zone
$n_{w}^{\sigma}$ number of workers in operating period $\sigma$
$n_{r}^{\sigma}$ number of robots in operating period $\sigma$
Appendix B: Proofs
B.1 Proof of Proposition 1
(i):
By definition 1, a feasible path of RC-S connects two active loading stations. If a feasible path has no more than two turns, all its turning points are within the blue region; otherwise, its endpoint will fall into a non-active loading station. The movement trajectories of occupied VPs will not cross the boundaries of the blue region; therefore, their upper limit is $\kappa$. Condition (i) is sufficient.
(ii):
We derive the average road length that each loading station can be allocated, noted as $\tilde{d}_{ls}$:
$\tilde{d}_{ls}(\alpha)$ $=$ $\frac{\big{(}1-{(1-\alpha)}^{2}\big{)}\cdot n_{h}n_{v}}{\frac{\alpha(n_{h}+n_{v})}{2}}$
$=$ $\frac{(2-\alpha)\cdot 2n_{h}n_{v}}{n_{h}+n_{v}}\geq(2-\alpha)\cdot min\{n_{h},n_{v}\}$
When the average travel distance of a task is less than $\tilde{d}_{ls}$, the proportion of the union of segments passed by the occupied VPs in the total set of segments is less than $\kappa$. Condition (ii) is sufficient.
$\square$
B.2 Derivation of Maximum Travel Distance in Constraint 5
We first examine the acceleration process of the robot after a turn. It needs to cover a distance of $2D$ within $2\tau_{e}$, reaching a final velocity of $v_{VP}$. We discuss two cases regarding the maximum travel distance the robot can cover while satisfying the final velocity $v_{VP}$, which are shown in Figure B-B-1 and B-B-1, respectively.
Figure B-1: Speed curves in the acceleration process
CASE 1: $2v_{max}-v_{VP}\geq 2\tau_{e}c_{max}$
The speed curve $v_{1}(t)$ and the associated travel distance $d_{1}$ are formulated as follows:
$v_{1}(t)=\left\{\begin{aligned} &c_{max}t,\quad t\leq\tau_{e}+\frac{v_{VP}}{2c_{max}}\\ &v_{VP}+c_{max}(2\tau_{e}-t),\quad\tau_{e}+\frac{v_{VP}}{2c_{max}}<t\leq 2\tau_{e}\\ \end{aligned}\right.$
$d_{1}=\int_{0}^{2\tau_{e}}v_{1}(t)=\frac{v_{VP}^{2}}{4c_{max}}+\tau_{e}v_{VP}+c_{max}\tau_{e}^{2}$
CASE 2: $2v_{max}-v_{VP}<2\tau_{e}c_{max}$
The speed curve $v_{1}(t)$ and the associated travel distance $d_{1}$ are formulated as follows:
$v_{2}(t)=\left\{\begin{aligned} &c_{max}t,\quad t\leq\frac{v_{max}}{c_{max}}\\ &v_{max},\quad\frac{v_{max}}{c_{max}}<t\leq 2\tau_{e}-\frac{v_{max}-v_{VP}}{c_{max}}\\ &v_{VP}+c_{max}(2\tau_{e}-t),\quad 2\tau_{e}-\frac{v_{max}-v_{VP}}{c_{max}}<t\leq 2\tau_{e}\\ \end{aligned}\right.$
$d_{2}=\int_{0}^{2\tau_{e}}v_{2}(t)=2\tau_{e}v_{max}-\frac{2v_{max}^{2}-2v_{max}v_{VP}+v_{VP}^{2}}{2c_{max}}$
Above all, the maximum travel distance is expressed by:
$d(\tau_{e},v_{VP},v_{max},c_{max})=\left\{\begin{aligned} &d_{1},\quad 2v_{max}-v_{VP}\geq 2\tau_{e}c_{max}\\ &d_{2},\quad 2v_{max}-v_{VP}<2\tau_{e}c_{max}\\ \end{aligned}\right.$
Similarly, the analysis for the acceleration process before a turn is consistent with that of the deceleration process, yielding the same result.
$\square$
B.3 Derivation of Average Travel Distance in Section 5
Let $I_{x}$ represents the set of x-coordinates of loading stations on the top and bottom sides, while $I_{y}$ represents the set of y-coordinates of loading stations on the left and right sides. Similarly, $W_{x}$ and $W_{y}$ denote the set of all coordinates of potential locations on the top and bottom sides, and on the left and right sides, respectively.
Area 1: Select turning points with equal probability in the central area and perform a weighted sum of all path lengths, with weights proportional to the route lengths.
$E[l_{1}]$ $=$ $\frac{2D}{\lvert I_{x}\rvert\cdot\lvert I_{y}\rvert}\cdot\sum_{x\in I_{x}}\sum_{y\in I_{y}}\{\frac{{(x+y)}^{2}}{2(n_{h}+n_{v})}+\frac{{[x+(n_{h}-y)]}^{2}}{2(n_{h}+n_{v})}+\frac{{[(n_{v}-x)+y]}^{2}}{2(n_{h}+n_{v})}$
$+\frac{{[(n_{v}-x)+(n_{h}-y)]}^{2}}{2(n_{h}+n_{v})}\}$
$=$ $2D\cdot[\frac{\frac{9+a^{2}}{6}(n_{h}^{2}+n_{v}^{2})-n_{h}n_{v}-\frac{1}{3}}{n_{h}+n_{v}}+1]$
Area 2: Calculate the average path lengths for starting points on the top and bottom sides and on the left and right sides, separately. According to the number of loading stations, perform a weighted sum. Based on the distance between the first turning point and the starting point, paths are divided into those that return to the original side and those that reach the opposite side, with each scenario having a probability of 1/2. For the former, the vertical movement distance is $n_{v}/2$ or $n_{h}/2$, while for the latter, it is $n_{v}$ or $n_{h}$, depending on the start point.
$E[l_{2}]$ $=$ $2D\cdot\big{\{}\frac{n_{v}}{n_{h}+n_{v}}\sum_{x\in I_{x}}\sum_{z\in I_{x}}\frac{\frac{1}{2}(\frac{n_{h}}{2}+\lvert z-x\rvert)+\frac{1}{2}(n_{h}+\lvert z-x\rvert)}{{\lvert I_{x}\rvert}^{2}}$
$+\frac{n_{h}}{n_{h}+n_{v}}\sum_{y\in I_{y}}\sum_{z\in I_{y}}\frac{\frac{1}{2}(\frac{n_{v}}{2}+\lvert z-y\rvert)+\frac{1}{2}(n_{v}+\lvert z-y\rvert)}{{\lvert I_{y}\rvert}^{2}}\big{\}}$
$=$ $2D\cdot[\frac{\alpha(n_{h}^{2}+n_{v}^{2})}{3(n_{h}+n_{v})}-\frac{2}{3\alpha(n_{h}+n_{v})}+\frac{3n_{h}n_{v}}{2(n_{h}+n_{v})}]$
Area 3: We first calculate the basic average path length. In accordance with workload balance, the starting and ending points of the path should be equally likely to fall on each workstation. Then add the detour lengths for the two different scenarios multiplied by their respective probabilities, according to the path allocation rules.
$E[l_{3}]$ $=$ $2D\cdot\sum_{x\in I_{x}}\sum_{y\in I_{y}}[\frac{x+y}{\lvert I_{x}\rvert\cdot\lvert I_{y}\rvert}+\frac{n_{h}}{2(n_{h}+n_{v})}\cdot\frac{\sum_{x\in W_{x}\backslash I_{x}}2(x-\frac{\alpha n_{v}}{2})}{\lvert W_{x}\backslash I_{x}\rvert}$
$+\frac{n_{v}}{2(n_{h}+n_{v})}\cdot\frac{\sum_{y\in W_{y}\backslash I_{y}}2(y-\frac{\alpha n_{h}}{2})}{\lvert W_{y}\backslash I_{y}\rvert}]$
$=$ $2D\cdot(\frac{n_{h}+n_{v}}{2}+\frac{(1+\alpha)}{4}\cdot\frac{n_{h}n_{v}}{n_{h}+n_{v}})$
B.4 Proof of Lemma 2
We first prove (i). By equation (27), the expression of average travel distance of a sorting task is as follows:
$\bar{l}(n_{h},n_{v},\alpha)=$ $2D\cdot[\frac{n_{h}^{2}+n_{v}^{2}}{n_{h}+n_{v}}\cdot\frac{-\alpha^{4}-5\alpha^{2}+18\alpha}{12}+\frac{n_{h}n_{v}}{n_{h}+n_{v}}\cdot\frac{-\alpha^{2}+2\alpha}{4}$
$+\frac{1}{n_{h}+n_{v}}\cdot\frac{3\alpha^{3}-\alpha^{2}-3\alpha-5}{12}+(n_{h}+n_{v})\cdot\frac{\alpha^{2}-2\alpha+1}{2}+\frac{-\alpha^{2}+2\alpha}{2}]$
where $\alpha$ is the ratio of the number of workers to the number of aisle entrances, $\alpha\in(0,1]$. We consider the partial derivative:
$\frac{\partial\bar{l}(n_{h},n_{v},\alpha)}{\partial\alpha}=$ $2D\cdot[\frac{n_{h}^{2}+n_{v}^{2}}{n_{h}+n_{v}}\cdot\frac{-4\alpha^{3}-10\alpha+18}{12}+\frac{n_{h}n_{v}}{n_{h}+n_{v}}\cdot\frac{-\alpha+1}{2}$
$+\frac{1}{n_{h}+n_{v}}\cdot\frac{9\alpha^{2}-2\alpha-3}{12}+(n_{h}+n_{v})(\alpha-1)-\alpha+1]$
At the two endpoints of the range of $\alpha$, the derivative has the value:
$\lim_{\alpha\xrightarrow{}0}\frac{\partial\bar{l}(n_{h},n_{v},\alpha)}{\partial\alpha}$ $=2D\cdot[\frac{3(n_{h}^{2}+n_{v}^{2})}{2(n_{h}+n_{v})}+\frac{n_{h}n_{v}}{2(n_{h}+n_{v})}-\frac{1}{4(n_{h}+n_{v})}-(n_{h}+n_{v})+1]$
$\leq 2D\cdot[\frac{n_{h}^{2}+n_{v}^{2}-3n_{h}n_{v}}{2(n_{h}+n_{v})}+1]<0$
$\leq 2D\cdot[-\frac{1}{4}min\{n_{h},n_{v}\}+1]\leq 0$
$\lim_{\alpha\xrightarrow{}1}\frac{\partial\bar{l}(n_{h},n_{v},\alpha)}{\partial\alpha}$ $=2D\cdot[\frac{n_{h}^{2}+n_{v}^{2}}{3(n_{h}+n_{v})}+\frac{1}{3(n_{h}+n_{v})}-1]$
$\geq 2D\cdot[\frac{1}{3}min\{n_{h},n_{v}\}-1]>0$
Similarly, we could calculate the second-order derivative and obtain that $\frac{\partial^{2}\bar{l}(n_{h},n_{v},\alpha)}{\partial\alpha^{2}}\geq 0$. Above all, we can prove that $\bar{l}(n_{h},n_{v},\alpha)$ decreases initially and then increases.
Next, we prove the advantage of square-shape site in (ii), namely the network with $n_{h}=n_{v}$. Without loss of generality, we assume $n_{h}=k\cdot n_{v}$, $k\geq 1$. Our objective is to show that the average travel distance is minimum when $k=1$. Let the area of the sorting zone be $S\cdot 4D^{2}$, then $n_{h}=\sqrt{kS}$, $n_{v}=\sqrt{S/k}$. Consider the derivative:
$\frac{\partial\bar{l}(S,k,\alpha)}{\partial k}=$ $2D\cdot[\frac{k^{2}-1}{2k^{\frac{3}{2}}}\cdot\sqrt{S}\cdot\frac{-\alpha^{4}-5\alpha^{2}+18\alpha}{6}+\frac{k-1}{2k^{\frac{1}{2}}{(k+1)}^{2}}\cdot\sqrt{S}\cdot\frac{-\alpha^{2}+2\alpha}{4}$
$+\frac{k-1}{2k^{\frac{1}{2}}{(k+1)}^{2}}\cdot\frac{1}{\sqrt{S}}\cdot\frac{3\alpha^{3}-\alpha^{2}-3\alpha-5}{12}+\frac{k-1}{2k^{\frac{3}{2}}}\cdot\sqrt{S}\cdot\frac{\alpha^{2}-2\alpha+1}{2}]$
It is easy to prove that the right side is consistently non-negative. As a result, the derivative is non-negative when $k\geq 1$, thus $\bar{l}(S,k,\alpha)\geq\bar{l}(S,1,\alpha)$.
To prove (iii), we first obtain the upper bound of $\bar{l}(n_{h},n_{v},\alpha)$. By (i), the upper bound can only be attained at the two endpoints:
$\bar{l}(n_{h},n_{v},0)=$ $n_{h}$
$\bar{l}(n_{h},n_{v},1)=$ $\frac{9}{8}n_{h}-\frac{1}{4n_{h}}+\frac{1}{2}<\frac{9}{8}n_{h}$
While $\alpha=0$ is not feasible for operation, we conclude $\bar{l}(n_{h},n_{v},\alpha)<\frac{9}{8}n_{h}$. According to the sorting demands with an average distribution, we can easily derive that the average distance for all three paths are greater than $n_{h}$, then we have: $\bar{l}(n_{h},n_{v},\alpha)<n_{h}$. The proof is completed.
$\square$
B.5 Proof of Proposition 2
We first obtain the expression of $\tilde{T}_{M}(n_{h},n_{v},n_{l})$:
$\tilde{T}_{M}(n_{h},n_{v},n_{l})=\frac{D}{\tau_{e}}\cdot\frac{n_{h}n_{v}}{a+b(n_{h}+n_{v})}\cdot\frac{\frac{2n_{l}}{n_{h}+n_{v}}-{(\frac{n_{l}}{n_{h}+n_{v}})}^{2}}{\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$
We further denote:
$G_{n_{h},n_{v}}(\alpha)=\frac{2\alpha-\alpha^{2}}{\bar{l}(n_{h},n_{v},\alpha)}$
$\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}=\frac{2(1-\alpha)\cdot\bar{l}(\cdot)-\frac{\partial\bar{l}(\cdot)}{\partial\alpha}\cdot(2\alpha-\alpha^{2})}{\bar{l}^{2}(\cdot)}$
From Lemma 1 and Lemma 2, we have:
$\lim_{\alpha\xrightarrow{}0}\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}=\frac{2}{\lim_{\alpha\xrightarrow{}0}\bar{l}(\cdot)}\geq\frac{2}{\sqrt{n_{h}n_{v}}}>0$
$\lim_{\alpha\xrightarrow{}1}\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}=\frac{-\frac{\partial\bar{l}(\cdot)}{\partial\alpha}}{\lim_{\alpha\xrightarrow{}0}\bar{l}^{2}(\cdot)}<0$
The non-negativity of the second-order derivative of $G_{n_{v},n_{h}}(\alpha)$ can be easily proved by obtaining the expression and checking the bounds of each term in the numerator. Hence, $G_{n_{v},n_{h}}(\alpha)$ has one zero point. The expression of the derivative of $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ can be written as:
$\frac{\partial\tilde{T}_{M}(n_{h},n_{v},n_{l})}{\partial n_{l}}=\frac{D}{\tau_{e}}\cdot\frac{n_{h}n_{v}}{a+b(n_{h}+n_{v})}\cdot\frac{1}{n_{h}+n_{v}}\cdot\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}$
It is immediate to prove that $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ has one zero point within the range $n_{l}\in(0,n_{h}+n_{v}]$ and it initially increases then decreases.
$\square$
Appendix C: Penalty Successive Linear Programming
The classic penalty successive linear programming (PSLP) enjoys good robustness and convergence properties for large-scale problems (Bazaraa et al.,, 2013). Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem along with the penalty function, and utilizes the concept of trust region (updated at each iteration) to control the step size. In each iteration $k$, a direction-finding linear program is formulated based on first-order Taylor series approximations to the objective and constraint functions, in addition to appropriate trust region restrictions on the direction components. The subproblem in iteration $k$ are formed as follows:
[…]
$s.t.\quad$ $y_{i}\geq g_{i}(\omega_{k})+\nabla{g_{i}(\omega_{k})}^{T}d,\quad i=1,2,\cdots,2\left|S\right|+2$ (C-1)
$-\Delta_{k}\leq d\leq\Delta_{k}$ (C-2)
$y_{i}\geq 0,\quad i=1,2,\cdots,2\left|S\right|+2$ (C-3)
where $\omega=\{n_{h},n_{v},n_{w}^{\sigma},n_{r}^{\sigma}\}$ is the set of all decision variables. $\Delta_{k}$ is the bound of $d$ in iteration $k$. $g_{i}(\omega_{k})$ represents the left-hand side of standard form inequality constraints (C-4)-(C-7) derived from the original LDP, assuming that the number of robots is less than the available VPs (considering the parking demand of robots during downtime, this setting is reasonable):
$-(n_{h}-1)(n_{v}-1)+N_{o}\leq 0$ (C-4)
[…]
$n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\qquad\forall\sigma\in\mathcal{S}$ (C-6)
$-\frac{D\cdot n_{r}^{\sigma}}{\tau_{e}\cdot E[l(n_{h},n_{v},\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})]}+{T}^{\sigma}\leq 0\qquad\forall\sigma\in\mathcal{S}$ (C-7)
[…]
Algorithm C-1 The PLSP algorithm to solve model LDP
0:    An initial feasible solution $\omega_{0}$onfidence intervals $0\leq\Delta_{LB}\leq\Delta_{1}$arameters $0\leq\rho_{0}\leq\rho_{1}\leq\rho_{2}\leq 1$, $\phi\leq 1$Large enough constant $\mu$Maximum iterations $N_{m}$
0:  The optimal value of layout decision variables, $\omega$
1:  Initialization: $k=1$, $R_{k}=0$, $iteration=1$
2:  while $iteration<N_{m}$ do
3:     repeat
[…]
5:        if $C_{EL_{k}}(\omega_{k})-C_{EL_{k}}(\omega_{k}+d_{k})=0$ or $\lceil\omega_{k}+d_{k}\rceil=\omega_{k}$ then
6:           Return $\omega_{k}$
7:           Stop
8:        else
9:           Calculate $R_{k}(\omega_{k},d_{k})$
10:        end if
11:        if $R_{k}<\rho_{0}$ then
12:           $\Delta_{k}=\alpha\Delta_{k}$
13:        end if
14:     until $R_{k}\geq\rho_{0}$
15:     $\omega_{k+1}=\lceil\omega_{k}+d_{k}\rceil$
16:     if $\rho_{0}\leq R_{k}<\rho_{1}$ then
17:        $\Delta_{k+1}=\phi\Delta_{k}$
18:     else if $\rho_{1}\leq R_{k}<\rho_{2}$ then
19:        $\Delta_{k+1}=\Delta_{k}$
20:     else if $R_{k}\geq\rho_{2}$ then
21:        $\Delta_{k+1}=\Delta_{k}/\phi$
22:     end if
23:     $\Delta_{k+1}=max\{\Delta_{k+1},\Delta_{LB}\}$
24:     $k=k+1$
25:  end while
In iteration $k$, the termination of algorithm could be determined by calculating the intermediate variables as follows:
$C_{E}(\omega_{k})=C_{d}(\omega_{k})+\mu\sum_{i}max\{0,g_{i}(\omega_{k})\}$ (C-8)
$C_{EL_{k}}(\omega_{k})=C_{d}(\omega_{k})+\nabla{C_{d}(\omega_{k})}^{T}d_{k}+\mu\sum_{i}max\{0,g_{i}(\omega_{k})+{\nabla g_{i}(\omega_{k})}^{T}\cdot d_{k}\}$ (C-9)
$R_{k}=\frac{C_{E}(\omega_{k})-C_{E}(\omega_{k}+d_{k})}{C_{EL_{k}}(\omega_{k})-C_{EL_{k}}(\omega_{k}+d_{k})}$ (C-10)
where $\mu$ is a large enough constant. Given that the variables in the original problem are integers, we introduced a rounding-up step for $\omega_{k}$ and the bounds in the search process. Additionally, due to the relatively limited solution space, the number of iterations in experiments remained below 50, and the total solving time was in the order of seconds. To ensure the feasibility of the solution for the original problem, the numerical update process includes a rounding-up step.